OpenStudy (anonymous):

From a point A on ground level, the angle of elevation to a plane is 75 degrees. From a point B on the ground, due west of A, the angle of elevation is 47 degrees. If A and B are 3500m apart, find the height of the plane off the ground.

OpenStudy (anonymous):

Start by drawing a diagram, similar to the one attached. Call G the point at the ground directly below P. Internal angle at A (a) = 180 - 75 = 105 degrees Internal angle at B (b) = 47 degrees Internal angle at P (p) = 180 - 105 - 47 = 28 degrees AB/sin(p) = AP/sin(b) 3500/sin(28) = AP/sin(47) AP = (3500/sin(28)) * sin(47) PG/sin(75) = AP/sin(90) PG = (AP/sin(90)) * sin(75) Make sense?

OpenStudy (anonymous):

okay that makes sense, Thankyou!

OpenStudy (anonymous):

No worries =)